OpenStudy (anonymous):

What is the inverse of f(x) = 2x − 8?

OpenStudy (anonymous):

To find the inverse of a function you are first gonna need to replace f(x) with y. Then you have to flip the x and y variables and solve for y. @sherlocked1270

OpenStudy (anonymous):

f(x) = 2x − 8 y = 2x − 8 x = 2y − 8 x+8=2y and so on..

OpenStudy (anonymous):

wait so what do I do after I get to x=2y+1

OpenStudy (johnkim2330):

you solve for y 2y=x-1 y=(x-1)/2

OpenStudy (usukidoll):

\(\color{blue}{\text{Originally Posted by}}\) @sunleaf01 f(x) = 2x − 8 y = 2x − 8 x = 2y − 8 x+8=2y and so on.. \(\color{blue}{\text{End of Quote}}\) After this step we are supposed to divide 2 throughout the equation.

OpenStudy (johnkim2330):

can someone help me with my question http://openstudy.com/updates/55a65b2be4b0f93dd7c4162d

OpenStudy (anonymous):

how do I remove the parentheses from Y=(x-1)/2

rishavraj (rishavraj):

how did u get tht??? y = 2x + 8 switch x and y x = 2y + 8 x - 8 = y x/2 - 4 = y now stick inverse notation

OpenStudy (usukidoll):

there's an error. Everything is fine until x+8=2y divide 2 throughout the entire equation to have y by itself the fraction part is correct, but we should have a 4

OpenStudy (anonymous):

ok I'm going on break see u later

rishavraj (rishavraj):

u will get x - 8 = 2y so (x - 8)/2 = y x/2 - 8/2 = y x/2 - 4 = y

OpenStudy (johnkim2330):

@rishavraj i ony saw sherlocked questin with x=2y+1 haha my bad i didnt see ur guys responses

OpenStudy (usukidoll):

that's alright. mistakes do happen ^^

OpenStudy (johnkim2330):

and can anyone help with this http://openstudy.com/updates/55a65b2be4b0f93dd7c4162d

OpenStudy (anonymous):

so u were saying that you have to get rid of the F(x) but in the answers that are provided It has f(x) f^-1(x)=1/2x-8 f^-1(x)=1/2x+4 f^-1(x)=1/2+8 f^-1(x)=2x+16

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

hi

rishavraj (rishavraj):

see my just previous reply... i ended with y = x/2 + 4 replace y from f^-1(x) so it woud be f^-1(x) = x/2 + 4

OpenStudy (anonymous):

oh ok thank u :D

rishavraj (rishavraj):

u welcome :))))

OpenStudy (anonymous):

ans is x+8 ------- 2